# Condensed Sets and the Solovay Model

### v1

@@ 3.2 $Sh(\mathbf{Sol}_{\neg\neg})\simeq\mathbf{VD}$
The definition of $\dot{x}_{F}$ makes sense for arbitrary presheaves $F$. The proofs of Lemmas 3.15 and 3.16 that $\Phi$ is fully faithful on maps $F\to G$ required only that $G$ be a sheaf. As a consequence, the assignment from a presheaf $F$ to the name $\dot{x}_{F}$ can be viewed as the composite $\Phi\circ\mathbf{a}$ with the sheafification functor. Explicitly, for a presheaf $F$ and name $\dot{a}$,
@@
Finally we are in a position to define the transfer functor, which is the left component of the geometric morphism defined between the previous two sections.
@@ 3.3.1 Weak Hausdorff spaces
Let $i\colon\mathbb{B}\to\mathbb{K}_{p}$ and $j\colon\mathbb{C}\to\mathbb{K}_{p}$ be the complete Boolean homomorphisms induced by $I,J$ respectively and let $\mathbb{D}\subset\mathbb{K}_{p}$ be the complete subalgebra generated by $\mathop{\textup{im}}(i)\cup\mathop{\textup{im}}(j)$. Then the computation in [6, Theorem 8.11], $f\circ S(i)=g\circ S(j)$ is the function which assigns $U\in S(\mathbb{D})$ to the unique point in
@@
Finally, we show that $\mathscr{F}$ is surjective. First, any maximal filter $\mathscr{F}$ on $K(X)^{V}$ which is in the Solovay model is definable from parameters added by some generic $I$ on a $\kappa$-small $\mathbb{B}$. By homogeneity of the tail forcing over $V[I]$, $\mathscr{F}\in V[I]$. Let $\dot{F}$ be a $\mathbb{B}$-name for a maximal filter on $K(X)^{V}$ such that $\dot{F}(I)=\mathscr{F}$ and fix $p\in\mathbb{B}$ such that for some $C\in K(X)^{V}$, $p$ forces $C\in\mathscr{F}$. Since $C$ is Hausdorff in the subspace topology, [6, Theorem 8.11] yields the desired continuous function.
@@ 3.3.2 Compact Hausdorff Spaces
As natural as Corollaries 3.25 and 3.26 may seem, they have some consequences that feel less natural. The main cause of this is that all atomless Boolean algebras of size $\kappa$ in $V$ become isomorphic in $V(\mathbb{R})$, with Stone spaces homeomorphic to the Cantor space. Although the homeomorphism is sometimes not in $\mathbf{VD}$, the Stone spaces consequently lose many topological properties.
In certain cases, there are other ways to map a compact space into a forcing extension that do not turn all Stone spaces into the Cantor space; we point interested readers to [48], where Todorčević develops machinery for mapping Rosenthal compacta into forcing extensions which behaves more naturally for this class of spaces. See also [36, Section 12] for another account of these results. We compute the images of the spaces from [36, Example 12.1] under $\Lambda$ to demonstrate the difference.
@@ 3.3.3 Large Baer-Specker Groups
In the category $\mathbf{VD}$, there is an isomorphism
$\Lambda\left(\prod_{I}\underline{\mathbb{Z}}\right)\cong\prod^{bdd}_{I}\mathbb{Z}.$
@@ 4.2 Putting it All Together
It must be stressed that aside from Lemmas 4.2 and 4.4, which are properties specific to the Solovay model, the above proof quite closely replicates exactly the same manipulations that lead to the freeness of $A$ in the proof due to Clausen–Scholze [14, Session 8].
@@
Corollary 4.8 (Bergfalk–Lambie-Hanson–Šaroch[6, Corollary 11.4]).
@@ 5 Extensions in $\mathbf{VD}$
The internal semantics can be used to formalise that the above suffices for the vanishing of $R\dashuline{Hom}^{1}(\mathbb{R},\mathbb{Z})$ in $\mathbf{VD}$. We will sketch why this is the case (an interested reader may find a more thorough exposition in the Stacks Project Chapter 21 [45, Tag 01FQ], more specifically [45, Tag 08J7]). Note that $\mathbf{VD}$ is a Grothendieck topos, hence its category of abelian groups permits $K$-injective resolutions (see [45, Tag 079P]). In turn this means that the internal $R\dashuline{Hom}^{1}(\mathbb{R},\mathbb{Z})$ can be computed by taking internal homs, i.e. it is the homology of the sequence
@@ 5.1 A New Choice Principle for the Solovay Model
Note that crucial in the above proof is the fact that $A_{r}$ is contained in $V(\mathbb{R})$, and in particular contains some elements which are definable from reals. In general, whether a countable $V$-definable set in $V^{\mathbb{K}}$ must have a $V$-definable element is open. The ordinal-definability version of this question was investigated by Kanovei–Lyubetsky [26, Theorem 1.2(ii)].
@@ 5.2 The Splitting
so that the cocycle defined from $\pi^{\prime}$ is also $\varphi$ a.e.. Strict here means that the inclusion of $\mathbb{Z}\hookrightarrow A^{\prime}$ is a closed embedding and $\pi^{\prime}$ is an open surjection. However, in this setting, $\mathbb{R}$ is projective (see for instance Moskowitz [37, Theorem 3.3]), hence $\pi^{\prime}$ has a continuous splitting, hence $\varphi$ is equal to a coboundary a.e., hence $\pi$ has a splitting as well. The coboundary can be chosen to be Borel over $V[r_{0}]$, hence it and the splitting live also in $V(\mathbb{R})$. Thus we have a splitting in $V(\mathbb{R})$.
@@ 6 Concluding Remarks
Let $\mathbf{Sol}_{\infty}$ and $\mathbf{Sol}_{\infty,\neg\neg}$ be obtained from $\mathbf{Sol}$ and $\mathbf{Sol}_{\neg\neg}$ by changing $({<}\kappa)$ to “set-sized”, and $\mathbf{VD}_{\infty}$ be the category of $V$-definable maps in the forcing extension by the class collapse forcing $Coll(\omega,{<}Ord)$. Is there a similar sequence of exact cocomplete functors
@@
If $\kappa$ is supercompact, does the elementary subtopos of $\mathbf{Pyk}_{\kappa}$ given by sheaves of size $\kappa$ permit a cocomplete Heyting functor to/from $\mathbf{Cond}$?
In another direction, Clausen–Scholze more recently turn to the topos of “light” condensed sets [12], constructed on the site of Stone spaces of countable weight. Several of the proofs in this paper adapt to the light context, though it is not clear to us if $Lan_{i}$ still preserves products. One must also be careful, as in Remark 3.20, to use $V(\mathbb{R}^{*})$ (the hull over the set of reals added in some intermediate extension).
@@ A Dealing in Absolutes
Uniqueness is clear, since two continuous maps into a Hausdorff space must agree on a closed set. For existence, recall the following criterion for $h$ to extend to $S$ (due originally to Taĭmanov [47]; see also Engelking [17, Theorem 3.2.1]): we must show that for $C_{1}$, $C_{2}$ disjoint closed in $X$, $h^{-1}C_{1}$, $h^{-1}C_{2}$ have disjoint closures in $S$.
@@
We now turn to the facts concerning absolutes and quasi-open maps. We refer the interested reader to Ponomarev–Shapiro [40] for more on this topic. Gleason shows that the spaces $\mathscr{A}(X)$ are projective in the category of compact Hausdorff spaces, hence every map possesses at least one absolute. Next we check that $\pi_{X}\colon\mathscr{A}(X)\to X$ is quasi-open (surjectivity and irreducibility being standard, due to Gleason [21]; see also the Stacks Project [45, Tag 08YH]).
@@ B The Who’s Who of $Sh(\mathbf{Sol})$
The natural map $i^{*}(A^{B})\to i^{*}A^{i^{*}B}$ between internal exponentials is a monomorphism. When $A\simeq Ran_{i}A^{\prime}$ or $B\simeq Lan_{i}B^{\prime}$, the map is a split monomorphism.
@@
We check pointwise:
$Sh(\mathbf{Sol})(C,i^{*}A^{i^{*}B})\simeq Sh(\mathbf{Sol})(C\times i^{*}B,i^{*}A)\simeq\mathbf{Pyk}(Lan_{i}C\times Lan_{i}i^{*}B,A)$
since $Lan_{i}$ preserves products, whereas
$Sh(\mathbf{Sol})(C,i^{*}(A^{B}))\simeq\mathbf{Pyk}(Lan_{i}C,A^{B})\simeq\mathbf{Pyk}(Lan_{i}C\times B,A).$
with the map between them induced by precomposition with the counit of the adjunction $Lan_{i}i^{*}B\to B$. Since $i^{*}$ is faithful, this counit is epi, hence precomposition with it induces a mono. If $B\simeq Lan_{i}B^{\prime}$, we have by the adjointness of $i^{*}$ and $Lan_{i}$, the retract
$Lan_{i}B^{\prime}\xrightarrow{Lan_{i}(\eta_{B^{\prime}})}Lan_{i}i^{*}Lan_{i}B^{\prime}\xrightarrow{\epsilon_{Lan_{i}B^{\prime}}}Lan_{i}B^{\prime}$
which gives the splitting.
Now if $A\simeq Ran_{i}A^{\prime}$, $Sh(\mathbf{Sol})(C,i^{*}A^{i^{*}B})$ above is isomorphic to
$Sh(\mathbf{Sol})(i^{*}Lan_{i}C\times i^{*}Lan_{i}i^{*}B,A^{\prime})$
and $\mathbf{Pyk}(Lan_{i}C\times B,A)$ is isomorphic to
$Sh(\mathbf{Sol})(i^{*}Lan_{i}C\times i^{*}B,A^{\prime})$
hence if we use the other retract
$i^{*}B\xrightarrow{\eta_{i^{*}B}}i^{*}Lan_{i}i^{*}B\xrightarrow{i^{*}(\epsilon_{B})}i^{*}B$
we have the required splitting. ∎
We also compute the subobject classifier of $\mathbf{Sol}$ and $\mathbf{Sol}_{\neg\neg}$. Recall (or see [32, Section III.7]) that the subobject classifier of $Sh(\mathbf{Sol})$ is given by
@@ I $I\mapsto\text{ sieve generated by }\{S(\mathbb{B}_{p})\hookrightarrow S(\mathbb{B})\mid p\in$ I\}
From this the verification that $Sh(\mathbf{Sol}_{\neg\neg})$ is the double negation topos of $Sh(\mathbf{Sol})$ is standard (one is localically embedded in the other, and $\Omega_{\mathbf{Sol}_{\neg\neg}}$ is the double-negation-fixed sublocale of $\Omega_{\mathbf{Sol}}$). Since we know that $Sh(\mathbf{Sol}_{\neg\neg})\simeq\mathbf{VD}$, the above can be translated to give the subobject classifier of $\mathbf{VD}$. For completeness we record it here.

### v2

@@ 3.2 $Sh(\mathbf{Sol}_{\neg\neg})\simeq\mathbf{VD}$
The definition of $\dot{x}_{F}$ makes sense for arbitrary presheaves $F$. The proofs of Lemmas 3.13 and 3.14 that $\Phi$ is fully faithful on maps $F\to G$ required only that $G$ be a sheaf. As a consequence, the assignment from a presheaf $F$ to the name $\dot{x}_{F}$ can be viewed as the composite $\Phi\circ\mathbf{a}$ with the sheafification functor. Explicitly, for a presheaf $F$ and name $\dot{a}$,
@@
Finally we are in a position to define the transfer functor, which is the left component of the geometric morphism defined by the previous two sections.
@@ 3.3.1 Weak Hausdorff spaces
Let $i\colon\mathbb{B}\to\mathbb{K}_{p}$ and $j\colon\mathbb{C}\to\mathbb{K}_{p}$ be the complete Boolean homomorphisms induced by $I,J$ respectively and let $\mathbb{D}\subset\mathbb{K}_{p}$ be the complete subalgebra generated by $\mathop{\textup{im}}(i)\cup\mathop{\textup{im}}(j)$. Then the computation in [5, Theorem 8.11], $f\circ S(i)=g\circ S(j)$ is the function which assigns $U\in S(\mathbb{D})$ to the unique point in
@@
Finally, we show that $\mathscr{F}$ is surjective. First, any maximal filter $\mathscr{F}$ on $K(X)^{V}$ which is in the Solovay model is definable from parameters added by some generic $I$ on a $\kappa$-small $\mathbb{B}$. By homogeneity of the tail forcing over $V[I]$, $\mathscr{F}\in V[I]$. Let $\dot{F}$ be a $\mathbb{B}$-name for a maximal filter on $K(X)^{V}$ such that $\dot{F}(I)=\mathscr{F}$ and fix $p\in\mathbb{B}$ such that for some $C\in K(X)^{V}$, $p$ forces $C\in\mathscr{F}$. Since $C$ is Hausdorff in the subspace topology, [5, Theorem 8.11] yields the desired continuous function.
@@ 3.3.2 Compact Hausdorff Spaces
As natural as Corollaries 3.23 and 3.24 may seem, they have some consequences that feel less natural. The main cause of this is that all atomless Boolean algebras of size $\kappa$ in $V$ become isomorphic in $V(\mathbb{R})$, with Stone spaces homeomorphic to the Cantor space. Although the homeomorphism is sometimes not in $\mathbf{VD}$, the Stone spaces consequently lose many topological properties.
In certain cases, there are other ways to map a compact space into a forcing extension that do not turn all Stone spaces into the Cantor space; we point interested readers to [46], where Todorcevic develops machinery for mapping Rosenthal compacta into forcing extensions which behaves more naturally for this class of spaces. See also [34, Section 12] for another account of these results. We compute the images of the spaces from [34, Example 12.1] under $\Lambda$ to demonstrate the difference.
@@ 3.3.3 Large Baer-Specker Groups
In the category $\mathbf{VD}$, there is an isomorphism $\Lambda\left(\prod_{I}\underline{\mathbb{Z}}\right)\cong\prod^{bdd}_{I}\mathbb{Z}.$
@@ 4.2 Putting it All Together
It must be stressed that aside from Lemmas 4.2 and 4.4, which are properties specific to the Solovay model, the above proof quite closely replicates exactly the same manipulations that lead to the freeness of $A$ in the proof by Clausen–Scholze [13, Session 8].
@@
Corollary 4.8 (Bergfalk–Lambie-Hanson–Šaroch[5, Corollary 11.4]).
@@ 5 Extensions in $\mathbf{VD}$
The internal semantics can be used to formalise that the above suffices for the vanishing of $R\dashuline{Hom}^{1}(\mathbb{R},\mathbb{Z})$ in $\mathbf{VD}$. We will sketch why this is the case (an interested reader may find a more thorough exposition in the Stacks Project Chapter 21 [43, Tag 01FQ], more specifically [43, Tag 08J7]). Note that $\mathbf{VD}$ is a Grothendieck topos, hence its category of abelian groups permits $K$-injective resolutions (see [43, Tag 079P]). In turn this means that the internal $R\dashuline{Hom}^{1}(\mathbb{R},\mathbb{Z})$ can be computed by taking internal homs, i.e. it is the homology of the sequence
@@ 5.1 A New Choice Principle for the Solovay Model
Note that crucial in the above proof is the fact that $A_{r}$ is contained in $V(\mathbb{R})$, and in particular contains some elements which are definable from reals. In general, whether a countable $V$-definable set in $V^{\mathbb{K}}$ must have a $V$-definable element is open. The ordinal-definability version of this question was investigated by Kanovei–Lyubetsky [24, Theorem 1.2(ii)].
@@ 5.2 The Splitting
so that the cocycle defined from $\pi^{\prime}$ is also $\varphi$ a.e.. Strict here means that the inclusion of $\mathbb{Z}\hookrightarrow A^{\prime}$ is a closed embedding and $\pi^{\prime}$ is an open surjection. However, in this setting, $\mathbb{R}$ is projective (see for instance Moskowitz [35, Theorem 3.3]), hence $\pi^{\prime}$ has a continuous splitting, hence $\varphi$ is equal to a coboundary a.e., hence $\pi$ has a splitting as well. The coboundary can be chosen to be Borel over $V[r_{0}]$, hence it and the splitting live also in $V(\mathbb{R})$. Thus we have a splitting in $V(\mathbb{R})$.
@@ 6 Concluding Remarks
Let $\mathbf{Sol}_{\infty}$ and $\mathbf{Sol}_{\infty,\neg\neg}$ be obtained from $\mathbf{Sol}$ and $\mathbf{Sol}_{\neg\neg}$ by changing $({<}\kappa)$ to “set-sized”, and $\mathbf{VD}_{\infty}$ be the category of $V$-definable maps in the forcing extension by the class collapse forcing $Coll(\omega,{<}Ord)$. Is there a similar sequence of exact cocontinuous functors
@@
If $\kappa$ is supercompact, does the elementary subtopos of $\mathbf{Pyk}_{\kappa}$ given by sheaves of size $\kappa$ permit a cocontinuous Heyting functor to/from $\mathbf{Cond}$?
In another direction, Clausen–Scholze more recently turn to the topos of “light” condensed sets [11], constructed on the site of Stone spaces of countable weight. Some of the proofs in this paper seem to adapt to the light context, though one must be careful to use the site of all countably generated complete Boolean algebras.
@@ A Dealing in Absolutes
Uniqueness is clear, since two continuous maps into a Hausdorff space must agree on a closed set. For existence, recall the following criterion for $h$ to extend to $S$ (due originally to Taĭmanov [45]; see also Engelking [16, Theorem 3.2.1]): we must show that for $C_{1}$, $C_{2}$ disjoint closed in $X$, $h^{-1}C_{1}$, $h^{-1}C_{2}$ have disjoint closures in $S$.
@@
We now turn to the facts concerning absolutes and quasi-open maps. We refer the interested reader to Ponomarev–Shapiro [38] for more on this topic. Gleason shows that the spaces $\mathscr{A}(X)$ are projective in the category of compact Hausdorff spaces, hence every map possesses at least one absolute. Next we check that $\pi_{X}\colon\mathscr{A}(X)\to X$ is quasi-open (surjectivity and irreducibility being standard, due to Gleason [19]; see also the Stacks Project [43, Tag 08YH]).
@@ B The Who’s Who of $Sh(\mathbf{Sol})$
Lemma B.3.
The natural map
$Lan_{i}(F\times G)\to Lan_{i}F\times Lan_{i}G$
is an objectwise surjection, hence in particular an epimorphism.
Proof.
The crucial point is, as before, that if $Y_{1}$, $Y_{2}$ are extremally disconnected, the maps $\pi_{i}:\mathscr{A}(Y_{1}\times Y_{2})\to Y_{i}$ are open. Next note that since products are computed pointwise
$Lan_{i}(F)\times Lan_{i}(G)=\{([u_{1}\in F(Y_{1}),f_{1}:X\to Y_{1}],[u_{2}\in G(Y_{2}),f_{2}:X\to Y_{2}])\}$
[…]
$([u_{1}\in F(Y_{1}),f_{1}:X\to Y_{1}],[u_{2}\in G(Y_{2}),f_{2}:X\to Y_{2}])\in Lan_{i}(F)\times Lan_{i}(G).$
Note that there is a unique $(f_{1},f_{2}):X\to Y_{1}\times Y_{2}$, and since $X$ is extremally disconnected, this lifts to a (not necessarily open) map $h:X\to\mathscr{A}(Y_{1}\times Y_{2})$. By construction, noting that the $\pi_{i}$ are open:
$u_{i}\in F(Y_{i}),f_{i}:X\to Y_{i}]=[u_{i}\in F(Y_{i}),\pi_{i}h]$
$F(\pi_{i})(u_{i})\in F(\mathscr{A}(Y_{1}\times Y_{2})),h:X\to\mathscr{A}(Y_{1}\times Y_{2})]$
However, this pair is now in the image of $Lan_{i}(F\times G)\to Lan_{i}(F)\times Lan_{i}(G)$ since
$(F(\pi_{1})(u_{1}),F(\pi_{2})(u_{2})),h]$
is a preimage. This shows that $Lan_{i}(F\times G)\to Lan_{i}(F)\times Lan_{i}(G)$ is an objectwise surjection. ∎
@@
The natural map $i^{*}(A^{B})\to i^{*}A^{i^{*}B}$ between internal exponentials is a monomorphism.
@@
We check objectwise:
$Sh(\mathbf{Sol})(C,i^{*}A^{i^{*}B})\simeq Sh(\mathbf{Sol})(C\times i^{*}B,i^{*}A)\simeq\mathbf{Pyk}(Lan_{i}(C\times i^{*}B),A)$
whereas
$Sh(\mathbf{Sol})(C,i^{*}(A^{B}))\simeq\mathbf{Pyk}(Lan_{i}C,A^{B})\simeq\mathbf{Pyk}(Lan_{i}C\times B,A),$
with the map between them given by precomposition with the map
$Lan_{i}(C\times i^{*}B)\xrightarrow{(Lan_{i}(\pi_{1}),Lan_{i}(\pi_{2}))}Lan_{i}C\times Lan_{i}i^{*}B\xrightarrow{(id,\epsilon_{B})}Lan_{i}C\times B.$
Since $i^{*}$ is faithful, $(id,\epsilon_{B})$ is epi, and by Lemma B.3, $(Lan_{i}(\pi_{1}),Lan_{i}(\pi_{2}))$ is an epi. Hence precomposition with this composite induces a mono. ∎
We also compute the subobject classifier of $\mathbf{Sol}$ and $\mathbf{Sol}_{\neg\neg}$. Recall (or see [29, Section III.7]) that the subobject classifier of $Sh(\mathbf{Sol})$ is given by
@@ I $I\mapsto\text{ sieve generated by }\{S(\mathbb{B}_{p})\hookrightarrow S(\mathbb{B})\mid p\in$ I\}
From this the verification that $Sh(\mathbf{Sol}_{\neg\neg})$ is the double negation topos of $Sh(\mathbf{Sol})$ is standard. Note first that a finite dense cover is a finite cover, hence $Sh(\mathbf{Sol}_{\neg\neg})$ is a full subcategory of $Sh(\mathbf{Sol})$. The double negation subtopos of $Sh(\mathbf{Sol})$ is the unique full subtopos whose subobject classifier is the double negation sublocale of $\Omega_{\mathbf{Sol}}$. The above proposition shows that $\Omega_{\mathbf{Sol}_{\neg\neg}}$ is the double-negation sublocale of $\Omega_{\mathbf{Sol}}$, which shows that $Sh(\mathbf{Sol}_{\neg\neg})$ is this double-negation subtopos.
Since we know that $Sh(\mathbf{Sol}_{\neg\neg})\simeq\mathbf{VD}$, the above can be translated to give the subobject classifier of $\mathbf{VD}$. For completeness we record it here.